OpenStudy (anonymous):

How do I identify the solution set for an inequality when given the graph: -x^2+9x-14(less than or equal to)0

OpenStudy (fibonaccichick666):

well, if given the graph, usually there is a shaded region

OpenStudy (anonymous):

I don't see a shaded region :(

OpenStudy (fibonaccichick666):

If you are given a graph, there has to be a shaded region for the graph to be correct for an inequality. But we can do it by hand and you can compare it to your graph if you wish

OpenStudy (fibonaccichick666):

oh oh oh wait a sec

OpenStudy (anonymous):

Yes, please...& ok

OpenStudy (fibonaccichick666):

I overcomplicated, sorry.

OpenStudy (fibonaccichick666):

so, you have something like this

OpenStudy (fibonaccichick666):

yea?

OpenStudy (anonymous):

On my paper it shows it on the right side but pretty much the same ......

OpenStudy (fibonaccichick666):

"-x^2+9x-14(less than or equal to)0" So when on that graph will you put an x value in and get a negative y value?

OpenStudy (fibonaccichick666):

or, a y value less than zero

OpenStudy (anonymous):

Below the x axis....

OpenStudy (fibonaccichick666):

yup, and at what x values does that occur?

OpenStudy (anonymous):

At 2 on the x axis and below. ... and at 7 and below

OpenStudy (fibonaccichick666):

for mine above, it looks like this the y value is less than zero on \[- \infty < x\le a ~~U~~b\le x<\infty\]